# Causal-Adapter: Taming Text-to-Image Diffusion for Faithful Counterfactual Generation

### v3

@@ Abstract
We present Causal-Adapter, a modular framework that adapts frozen text-to-image diffusion backbones for counterfactual image generation. Our method enables causal interventions on target attributes, consistently propagating their effects to causal dependents without altering the core identity of the image. In contrast to prior approaches that rely on prompt engineering without explicit causal structure, Causal-Adapter leverages structural causal modeling augmented with two attribute regularization strategies: prompt-aligned injection, which aligns causal attributes with textual embeddings for precise semantic control, and a conditioned token contrastive loss to disentangle attribute factors and reduce spurious correlations. Causal-Adapter achieves state-of-the-art performance on both synthetic and real-world datasets, with up to 91% MAE reduction on Pendulum for accurate attribute control and 87% FID reduction on ADNI for high-fidelity MRI image generation. These results show that our approach enables robust, generalizable counterfactual editing with faithful attribute modification and strong identity preservation.
@@ 1 Introduction
Answering counterfactual questions (e.g. inferring what an event would have happened under an alternative action) requires understanding the cause–effect relationships among variables and performing hypothetical reasoning (Pearl, 2009; Schölkopf et al., 2021; Weinberg et al., 2024).  Classical generative models typically tackle counterfactual-style tasks such as image editing or style transfer in a non-causal manner, whereas subsequent work augments such models with explicit structural causal models (SCMs) to implement abduction–action–prediction (Pearl, 2013; Pawlowski et al., 2020; De Sousa Ribeiro et al., 2023). This design drives advances in counterfactual image generation (Figure 1) by enforcing edits consistent with an implied causal graph which enables critical domain-specific counterfactual generation applications, such as simulating medical images with fine-grained anatomical changes associated with aging or disease progression to improve clinical interpretation (Starck et al., 2025). Faithful counterfactual generation remains challenging, as real-world attributes are often causally entangled (e.g., only males can grow beards). This entanglement complicates disentangling factors and generalizing edits, making it nontrivial to ensure that interventions yield the intended visual effect while keeping non-intervened attributes invariant and preserving identity-specific details (Komanduri et al., 2024a).
Counterfactual Image Generation. Early approaches modeled counterfactual image generation using normalizing flows (Papamakarios et al., 2021; Winkler et al., 2020), variational autoencoders (VAEs) (Kingma & Welling, 2013; Yang et al., 2021), hierarchical VAEs (Vahdat & Kautz, 2020; De Sousa Ribeiro et al., 2023), and generative adversarial networks (GANs) (Goodfellow et al., 2014; Kocaoglu et al., 2018), and encouraged attribute disentanglement through variational objectives (Higgins et al., 2017). However, variational optimization inevitably introduces uncertainty into the latent space, which can lead to posterior collapse of meaningful factors, creating a trade-off between image fidelity and controllable attribute manipulation (Figure 2a). Recent works integrate diffusion models with the SCM, capitalizing on its high perceptual quality to explore counterfactual identifiability (Sanchez & Tsaftaris, 2022; Rasal et al., 2025; Komanduri et al., 2024c; Pan & Bareinboim, 2024; Xia et al., 2025). Despite domain-specific tuning, previous approaches perform disentanglement only in auxiliary encoders, which has limited effect on diffusion latents and leads to incomplete disentanglement (Figure 2b and c). This makes models prone to spurious correlations, where target factor interventions often cause unintended changes in non-intervened attributes.
Text-to-Image based Editing. An alternative perspective is to view counterfactual image generation as a text-to-image (T2I) based editing, which typically relies on an inversion process (Song et al., 2021). The aim is to generate a target edited image from projected latent states with condition manipulation (Hertz et al., 2023; Ho & Salimans, 2021). To reduce reconstruction error and preserve essential contents, several methods optimize either the unconditional text embedding (Mokady et al., 2023; Xu et al., 2024; Miyake et al., 2025; Ju et al., 2024; Dong et al., 2023), or learn conditional concept embeddings for better attribute disentanglement (Gal et al., 2023; Vinker et al., 2023; Jin et al., 2024). However, generic T2I based editing remains insufficient for counterfactual generation. Existing methods heavily rely on carefully engineered inversion prompts to obtain reliable editing guidance. These approaches lack an explicit, learnable SCM over semantic attributes, making them difficult to guarantee both causal faithfulness and identity preserved counterfactual image generation (Figure 2d). A broader discussion of related works can be found in Appendix A.
@@
Herein, we propose Causal-Adapter, an adaptive and modular framework that tames text-to-image diffusion model, such as Stable-Diffusion (SD) (Rombach et al., 2022), for counterfactual generation. Unlike prior diffusion-based methods that require considerable re-training or fine-tuning  (Sanchez & Tsaftaris, 2022; Komanduri et al., 2024c; Pan & Bareinboim, 2024; Rasal et al., 2025), our method simply injects causal semantic attributes into a frozen backbone via a pluggable adapter (Figure 2e). Inspired by advances in controllable diffusion (Zhang et al., 2023; Zhao et al., 2023; Mou et al., 2024; Li et al., 2025), we investigate integrating semantic attributes with image embeddings as additional conditions. We find that naive fusion fails to achieve sufficient disentanglement or semantic alignment with spatial features in the diffusion latents. To address this, we introduce two regularization strategies that align causal semantic attributes with textual embeddings, improving disentanglement and causal faithfulness (Figure 2f). Our main contributions are summarized as follows:
@@
  • We propose Causal-Adapter, a framework that employs an adapter encoder to learn causal interactions between semantic variables. These interactions are injected into a frozen diffusion backbone and jointly optimized, introducing a dynamic prior that enables faithful counterfactual generation.
@@ 2 Methodology
Figure 3: Motivational study and preliminary counterfactual generation results between T2I methods and Causal-Adapter. (a) Fine-grained anatomical counterfactual editing of brain ventricular volume using inversion-based editing (NTI (Mokady et al., 2023)), multi-concept prompt-learning editing (MCPL (Jin et al., 2024)), and our approach. (b) Comparison of counterfactual editing results on human faces. (c) Averaged cross-attention maps from the base Causal-Adapter and the Causal-Adapter with regularizers. Full results and technical details are presented in Appendix B.
We first present the preliminaries in Section 2.1, followed by a motivational study in Section 2.2. This study examines the systematic limitations of current T2I based editing methods for counterfactual generation and emphasizes the need to leverage causal semantic attributes for producing faithful counterfactuals (Figure 3). Motivated by these findings, we further introduce the proposed Causal-Adapter in Section 2.3 and describe our regularization strategies in Section 2.4, which further enhance counterfactual generation through semantic disentanglement.
@@ 2.1 Preliminaries
Counterfactual reasoning aims to answer queries “Given observation $Y$, what would $y_{i}$ have been if its parents $\mathrm{Pa}_{i}$ had been different?” This is formalized by the abduction–action–prediction procedure (Pearl, 2013; Pawlowski et al., 2020; Shen et al., 2022):
  1. 1.
    Abduction: inferring posterior over exogenous variables consistent with observation, $p_{\mathcal{S}}(U\mid Y)$.
  2. 2.
@@
  3. 3.
[…]
Text-to-image Diffusion Model (T2I).
T2I models are a class of probabilistic generative models that synthesize images conditioned on textual prompts by progressively denoising Gaussian noise (Sohl-Dickstein et al., 2015; Ho et al., 2020; Rombach et al., 2022). Given an input image $x$, a frozen image encoder $\mathcal{E}$ maps it to a latent representation $z_{0}=\mathcal{E}(x)$. A conditional text embedding $V=c_{\phi}(\texttt{Prompt})$ is obtained from a pre-trained text encoder $c_{\phi}$ with parameters $\phi$, where Prompt denotes the input text. To train the denoising network $\epsilon_{\theta}$, Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to $z_{0}$ to obtain a noisy latent $z_{t}$ at timestep $t\sim\mathrm{Uniform}\{1,\dots,T\}$. The denoising loss $\mathcal{L}_{\text{DM}}$ is defined as:
$\min_{\theta,\phi}\;\mathbb{E}_{\mathcal{E}(x),V,\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,V)\right\|_{2}^{2}\right].$ (1)
@@ 2.3 Overview of Causal-Adapter
Recent advances on controllable diffusion methods (Zhang et al., 2023; Zhao et al., 2023; Mou et al., 2024; Li et al., 2025) show that a frozen T2I diffusion backbone can be steered by auxiliary control signals (e.g., segmentation masks or human poses) supplied through a trainable side module. We adopt the same high-level recipe by treating causal semantic attributes as the auxiliary control signals. The causal mechanism between attributes is then learned and injected explicitly into the diffusion backbone via a compact modular encoder that we term as Causal-Adapter (Figure 4).
@@
$\mathcal{L}_{\text{NLL}}\coloneqq-\sum_{i=1}^{K}\!\log p\bigl(y_{i}\mid f_{i}(\mathrm{Pa}_{i},u_{i})\bigr)\;=\;\frac{1}{2}\sum_{i=1}^{K}\frac{\lVert y_{i}-\bar{y}_{i}\rVert_{2}^{2}}{\sigma_{i}^{2}}+\log(2\pi\sigma_{i}^{2}).$ (3)
@@
Prior attempts have treated causal attributes as the conditioning signal to the diffusion backbone, which is computationally expensive as it requires retraining the entire model when switching across datasets (Sanchez & Tsaftaris, 2022; Komanduri et al., 2024c; Rasal et al., 2025). To overcome this limitation, we introduce a branch module that delivers causal guidance without modifying the backbone parameters to preserve the flexibility and scalability of a pretrained T2I model. Given the image latent representation $z_{t}=\mathcal{E}(x,t)$ at diffusion step $t$ and the text embeddings $V$, we instantiate a half-scale replica $\ddot{\epsilon}_{\psi}$ of the denoiser $\epsilon_{\theta}$. This replica specializes $\ddot{\epsilon}_{\psi}$ in injecting control signals and is parameterized by $\psi$. The causal attributes $Y$ are added to $z_{t}$ and processed by $\ddot{\epsilon}_{\psi}$:
@@
the residuals $r_{t}$ are injected into the mid- and up-sampling blocks of the frozen denoiser $\epsilon_{\theta}$. During training, we keep all backbone parameters fixed and optimize only $\psi$ under the loss in Eqn. 1.
Algorithm 1 Causal-Adapter: Training
1:Input: image $x$, semantic attributes $Y$, frozen $\{\mathcal{E},c_{\phi},\epsilon_{\theta}\}$, text embeddings $V$.
2:Output: trained causal mechanisms $F$ and adapter $\ddot{\epsilon}_{\psi}$.
3:# training causal mechanisms
4: Fit $F$ on $Y$ by minimising Eqn. 3
5:# training conditional adapter
6:for $step=1$ to $T$ do
7:  Encode $z_{t}=\mathcal{E}(x,t)$
8:  $r_{t}\coloneqq\ddot{\epsilon}_{\psi}(z_{t}\oplus Y,t,V)$
9:  $\psi:=\operatorname*{arg\,min}_{\psi}E_{z,r,V,\epsilon,t}[\|\epsilon-\epsilon_{\theta}(z_{t},r_{t},t,V)\|_{2}^{2}]$
10:end for
11:Return $F$ and $\ddot{\epsilon}_{\psi}$.
Algorithm 2 Causal-Adapter: Inference
1:Input: image $x$, semantic attributes $Y$, frozen $\{\mathcal{E},c_{\phi},\epsilon_{\theta}\}$, text embeddings $V$, trained $\{F,\ddot{\epsilon}_{\psi}\}$
2:Output: counterfactual image $\bar{x}$.
3:Encode $z_{0}=\mathcal{E}(x,0)$
4:for $t=0$ to T1T-1 do
5:  $r_{t}\coloneqq\ddot{\epsilon}_{\psi}(z_{t}\oplus Y,\,t,\,V)$
6:  $z_{t+1}^{\star}\coloneqq H_{\theta}(z_{t},\,r_{t},\,V,\,t)$// abduction
7:end for
[…]
9: Initialize $\bar{z}_{T}\coloneqq z_{T}^{\star}$
10:for $t=T$ down to 11 do
11:  $\bar{r}_{t}\coloneqq\ddot{\epsilon}_{\psi}(z_{t}^{\star}\oplus\bar{Y},\,t,\,V)$
12:  $\bar{z}_{t-1}\coloneqq H_{\theta}^{-1}(\bar{z}_{t},\,\bar{r}_{t},\,V,\,t)$// prediction
13:end for
14:Return $\bar{x}=\textrm{Decode}(\bar{z}_{0})$
@@
We implement counterfactual generation under the abduction–action–prediction procedure via DDIM inversion and sampling (Song et al., 2021). Let $H_{\theta}(z_{t},V,t)$ denote the DDIM inversion operator and let $H_{\theta}^{-1}$ be its generative inverse. The procedure is as follows: (1) Abduction: given an observed image $x$, we inject the original condition residuals $r_{t}$ and run $H_{\theta}(z_{t=0},r_{t},V,t)$ to recover the corresponding exogenous noise $z_{T}^{\star}$. (2) Action: we intervene $y_{i}$ by setting $y_{i}\!\leftarrow\!\bar{y_{i}}$ via $do(y_{i}=\bar{y_{i}})$, propagate the attribute change through Eqn. 2 and Eqn. 4 to obtain the updated $\bar{Y}=\{\bar{y_{i}}\}_{k=1}^{K}$ and its residuals $\bar{r_{t}}$. (3) Prediction: the counterfactual outcome is generated with $H_{\theta}^{-1}(z_{t=T}^{\star},\bar{r}_{t},V,t)$. Optionally, classifier-free guidance (CFG) (Ho & Salimans, 2021) can be applied to amplify the counterfactual signal with guidance weight $\alpha$:
@@ 2.4 Regularizing the Causal-Adapter
Causal conditions $Y$ are numeric attributes that lack direct alignment with pixel-level representations but are semantically closer to text embeddings. Inspired by prior finding (Yang et al., 2024) that disentangled tokens in the prompt can guide cross-attention to align semantics with spatial structure in diffusion latents, we inject $Y$ through the prompt channel. This allows the cross-attention module to propagate attribute semantics into spatial image features during generation. We introduce a prompt-aligned injection (PAI) mechanism that maps each causal attribute to a learnable token embedding. Let $C=[\,c_{1},\ldots,c_{K}\,]^{\top}\in\mathbb{R}^{K\times d}$ denote the embeddings of $K$ placeholder tokens with dimension $d$. Each attribute $y_{i}$ is mapped into the text space via a small linear projector $g_{i}:\mathbb{R}\!\to\!\mathbb{R}^{d}$. We form the attribute-injected token embeddings:
$v_{i}(y_{i})\;=\;c_{i}+g_{i}(y_{i}),\qquad V(Y)\;=\;[\,v_{1}(y_{1}),\ldots,v_{K}(y_{K})\,]^{\top}\in\mathbb{R}^{K\times d}.$ (6)
The set $V(Y)$ serve as the conditioning input to both the frozen denoiser and the adapter via cross-attention. During training, we jointly optimize the adapter $\psi$, the placeholders token embeddings $C$, and the projectors $G=\{g_{i}\}_{i=1}^{K}$ using the standard MSE:
@@
$\mathcal{L}_{\text{CTC}}=\frac{1}{BK}\sum_{k=1}^{K}\sum_{b=1}^{B}\left[-\,\operatorname{sim}(v_{b}^{k},v_{b^{\prime}}^{k})/\tau+\log\!\Big(\sum_{\begin{subarray}{c}k^{\prime}=1,k^{\prime}\neq k\end{subarray}}^{K}\sum_{\begin{subarray}{c}b^{\prime}=1,b^{\prime}\neq b\end{subarray}}^{B}\exp\!\bigl(\operatorname{sim}(v_{b}^{k},v_{b^{\prime}}^{k^{\prime}})/\tau\bigr)\Big)\right]$ (8)
@@
$\lambda$ denotes a scaling coefficient. The proposed regularizers improve semantic–spatial alignment in the latent space (Figure 3c) and reduce spurious correlations (Figure 9c). The learned attention maps can be leveraged for localized editing through attention-guided manipulation. Following  Ju et al. (2024), interventions are applied only to cross-attention weights corresponding to targeted tokens, while attention for non-intervened tokens is preserved. This ensures that edits remain spatially localized (e.g., gender) without altering unrelated attributes (e.g., hairstyle), improving identity preservation without introducing extra training objectives. Full algorithm is presented in Appendix. C.3.

### v5

@@ Abstract
We present Causal-Adapter, a modular framework that adapts frozen text-to-image diffusion for counterfactual generation. Our method enables causal interventions on target attributes, consistently propagating their effects to causal dependents without altering the core identity of the image. In contrast to prior approaches that rely on prompt engineering without explicit causal mechanism, Causal-Adapter leverages structural causal modeling augmented with two attribute regularization strategies: prompt-aligned injection, which aligns causal attributes with textual embeddings for precise semantic control, and a conditioned token contrastive loss to disentangle attribute factors and reduce spurious correlations. Causal-Adapter achieves state-of-the-art performance on both synthetic and real-world datasets, with up to 91% MAE reduction on Pendulum for accurate attribute control and 87% FID reduction on ADNI for high-fidelity MRI generation. These results show that our approach enables robust, generalizable counterfactual editing with faithful attribute modification and strong identity preservation.
Machine Learning, ICML
@@ 1 Introduction
Answering counterfactual questions (e.g. inferring what an event would have happened under an alternative action) requires understanding the cause–effect relationships among variables and performing hypothetical reasoning (Pearl, 2009; Schölkopf et al., 2021; Weinberg et al., 2024).  Classical generative models typically tackle counterfactual-style tasks such as image editing or style transfer in a non-causal perspective. Taking a pre-defined causal graph and semantic attributes as standard inputs, subsequent work augments such models with explicit structural causal models (SCMs) to implement abduction–action–prediction (Pearl, 2013; Pawlowski et al., 2020; De Sousa Ribeiro et al., 2023; Komanduri et al., 2024c; Rasal et al., 2025). This design drives advances in counterfactual image generation (Figure 1) by enforcing edits consistent with a causal graph which enables critical domain-specific counterfactual generation applications, such as simulating medical images with fine-grained anatomical changes associated with aging or disease progression to improve clinical interpretation (Starck et al., 2025). Faithful counterfactual generation remains challenging, as real-world attributes are often causally entangled (e.g., only males can grow beards). This entanglement complicates disentangling factors and generalizing edits, making it nontrivial to ensure that interventions yield the intended visual effect while keeping non-intervened attributes invariant and preserving identity-specific details (Komanduri et al., 2024a).
@@
Counterfactual Image Generation. Early approaches modeled counterfactual image generation using normalizing flows (Papamakarios et al., 2021; Winkler et al., 2020), variational autoencoders (VAEs) (Kingma and Welling, 2013; Yang et al., 2021), hierarchical VAEs (Vahdat and Kautz, 2020; De Sousa Ribeiro et al., 2023), and generative adversarial networks (GANs) (Goodfellow et al., 2014; Kocaoglu et al., 2018), and encouraged attribute disentanglement through variational objectives (Higgins et al., 2017). However, variational optimization inevitably introduces uncertainty into the latent space, which can lead to posterior collapse of meaningful factors, creating a trade-off between image fidelity and controllable attribute manipulation (Figure 2a). Recent works integrate diffusion models with the SCM, capitalizing on its high perceptual quality to explore counterfactual identifiability (Sanchez and Tsaftaris, 2022; Rasal et al., 2025; Komanduri et al., 2024c; Pan and Bareinboim, 2024; Xia et al., 2025). Despite domain-specific tuning, previous approaches perform disentanglement only in auxiliary encoders, which has limited effect on diffusion latents and leads to incomplete disentanglement (Figure 2b and c). This makes models prone to spurious correlations, where target factor interventions often cause unintended changes in non-intervened attributes.
Text-to-Image based Editing. An alternative perspective is to view counterfactual image generation as a text-to-image (T2I) based editing, which typically relies on an inversion process (Song et al., 2021). The aim is to generate a target edited image from projected latent states with condition manipulation (Hertz et al., 2023; Ho and Salimans, 2021). To reduce reconstruction error and preserve essential contents, several methods optimize either the unconditional text embedding (Mokady et al., 2023; Xu et al., 2024; Miyake et al., 2025; Ju et al., 2024; Dong et al., 2023), or learn conditional concept embeddings for better attribute disentanglement (Gal et al., 2023; Vinker et al., 2023; Jin et al., 2024). However, generic T2I based editing remains insufficient for counterfactual generation. Existing methods heavily rely on carefully engineered inversion prompts to obtain reliable editing guidance. These approaches lack an explicit, learnable SCM over semantic attributes, making them difficult to guarantee both causal faithfulness and identity preserved counterfactual image generation (Figure 2d). A broader discussion of related works can be found in Appendix A.
Figure 3: Motivational study and preliminary counterfactual generation results between T2I methods and Causal-Adapter. (a) Fine-grained anatomical counterfactual editing of brain ventricular volume using inversion-based editing (NTI (Mokady et al., 2023)), multi-concept prompt-learning editing (MCPL (Jin et al., 2024)), and our approach. (b) Comparison of counterfactual editing results on human faces. (c) Cross-attention maps: base vs. regularized Causal-Adapter. Details are presented in Appendix B.
Herein, we propose Causal-Adapter, an adaptive and modular framework that tames text-to-image diffusion model, such as Stable-Diffusion (SD) (Rombach et al., 2022), for counterfactual generation. Unlike prior diffusion-based methods that require considerable re-training or fine-tuning  (Sanchez and Tsaftaris, 2022; Komanduri et al., 2024c; Pan and Bareinboim, 2024; Rasal et al., 2025), our method simply injects causal semantic attributes into a frozen backbone via a pluggable adapter (Figure 2e). Inspired by advances in controllable diffusion (Zhang et al., 2023; Zhao et al., 2023; Mou et al., 2024; Li et al., 2025), we investigate integrating semantic attributes with image embeddings as additional conditions. We find that naive fusion fails to achieve sufficient disentanglement or semantic alignment with spatial features in the diffusion latents. To address this, we introduce two regularization strategies that align causal semantic attributes with textual embeddings, improving disentanglement and causal faithfulness (Figure 2f). Our main contributions are summarized as follows:
@@
  • We propose Causal-Adapter, a framework that employs an adapter encoder to learn causal interactions between semantic variables. These interactions are injected into a frozen diffusion and jointly optimized, introducing a dynamic prior for faithful counterfactual generation.
@@ 2.3 Overview of Causal-Adapter
Recent advances on controllable diffusion methods (Zhang et al., 2023; Mou et al., 2024; Li et al., 2025) show that a frozen T2I diffusion backbone can be steered by auxiliary control signals (e.g., segmentation masks or human poses) supplied through a trainable side module. We adopt the same high-level recipe by treating causal semantic attributes as the auxiliary control signals. The causal mechanism between attributes is then learned and injected explicitly into the diffusion backbone via a compact modular encoder that we term as Causal-Adapter (Figure 4).
@@
$\mathcal{L}_{\text{NLL}}$ $=-\sum_{i=1}^{K}\!\log p\bigl(y_{i}\mid f_{i}(\mathrm{Pa}_{i},u_{i})\bigr)$ (2)
$=\;\frac{1}{2}\sum_{i=1}^{K}\frac{\lVert y_{i}-\bar{y}_{i}\rVert_{2}^{2}}{\sigma_{i}^{2}}\;+\;\log(2\pi\sigma_{i}^{2}).$
@@
Prior attempts have treated causal attributes as the conditioning signal to the diffusion backbone, which is computationally expensive as it requires retraining the entire model when switching across datasets (Sanchez and Tsaftaris, 2022; Komanduri et al., 2024c; Rasal et al., 2025). To overcome this limitation, we introduce a branch module that delivers causal guidance without modifying the backbone parameters to preserve the flexibility and scalability of a pretrained T2I model. Given the image latent representation $z_{t}=\mathcal{E}(x,t)$ at diffusion step $t$ and the text embeddings $V$, we instantiate a half-scale replica $\ddot{\epsilon}_{\psi}$ of the denoiser $\epsilon_{\theta}$. This replica specializes $\ddot{\epsilon}_{\psi}$ in injecting control signals and is parameterized by $\psi$. The causal attributes $Y$ are added to $z_{t}$ and processed by $\ddot{\epsilon}_{\psi}$:
@@
the residuals $r_{t}$ are injected into the mid- and up-sampling blocks of the frozen denoiser $\epsilon_{\theta}$. During training, we keep all backbone parameters fixed and optimize only $\psi$ with a T2I noise prediction objective.
@@
We implement counterfactual generation under the abduction–action–prediction procedure via DDIM inversion and sampling (Song et al., 2021). Let $H_{\theta}(z_{t},V,t)$ denote the DDIM inversion operator and let $H_{\theta}^{-1}$ be its generative inverse. The procedure is as follows: (1) Abduction: given an observed image $x$, we inject the condition residuals $r_{t}$ and run $H_{\theta}(z_{t=0},r_{t},V,t)$ to recover the corresponding exogenous noise $z_{T}^{\star}$. (2) Action: we intervene $y_{i}$ by setting $y_{i}\!\leftarrow\!\bar{y_{i}}$ via $do(y_{i}=\bar{y_{i}})$, propagate the attribute change through Eqn. 1 and Eqn. 3 to obtain the updated $\bar{Y}=\{\bar{y_{i}}\}_{k=1}^{K}$ and its residuals $\bar{r_{t}}$. (3) Prediction: the counterfactual outcome is generated with $H_{\theta}^{-1}(z_{t=T}^{\star},\bar{r}_{t},V,t)$. Optionally, classifier-free guidance (CFG) (Ho and Salimans, 2021) can be applied to amplify the counterfactual signal with guidance weight $\alpha$:
@@ 2.4 Regularizing the Causal-Adapter
Causal conditions $Y$ are numeric attributes that lack direct alignment with pixel-level representations but are semantically closer to text embeddings. Inspired by prior finding (Yang et al., 2024) that disentangled tokens in the prompt can guide cross-attention to align semantics with spatial structure in diffusion latents, we inject $Y$ through the prompt channel. This allows the cross-attention module to propagate attribute semantics into spatial image features during generation. We introduce a prompt-aligned injection (PAI) mechanism that maps each causal attribute to a learnable token embedding. Let $C=[\,c_{1},\ldots,c_{K}\,]^{\top}\in\mathbb{R}^{K\times d}$ denote the embeddings of $K$ placeholder tokens with dimension $d$. Each attribute $y_{i}$ is mapped into the text space via a linear projector $g_{i}:\mathbb{R}\!\to\!\mathbb{R}^{d}$. We form the attribute-injected token embeddings:
$v_{i}(y_{i})$ $=c_{i}+g_{i}(y_{i}),\qquad i=1,\ldots,K,$ (5)
$V(Y)$ $=\bigl[\,v_{1}(y_{1}),\ldots,v_{K}(y_{K})\,\bigr]^{\top}\in\mathbb{R}^{K\times d}.$
The set $V(Y)$ serve as the conditioning input to both the frozen denoiser and the adapter via cross-attention. During training, we jointly optimize the adapter $\psi$, the placeholders token embeddings $C$, and the projectors $G=\{g_{i}\}_{i=1}^{K}$ using the standard noise prediction loss $\mathcal{L}_{\text{DM}}$:
@@
$\mathcal{L}_{\text{CTC}}=\frac{1}{BK}\sum_{k=1}^{K}\sum_{b=1}^{B}\Bigg[-\frac{\operatorname{sim}\!\left(v_{b}^{k},\,v_{b^{\prime}}^{k}\right)}{\tau}+\log\!\Bigg(\sum_{\begin{subarray}{c}k^{\prime}=1\\ k^{\prime}\neq k\end{subarray}}^{K}\sum_{\begin{subarray}{c}b^{\prime}=1\\ b^{\prime}\neq b\end{subarray}}^{B}\exp\!\left(\frac{\operatorname{sim}\!\left(v_{b}^{k},\,v_{b^{\prime}}^{k^{\prime}}\right)}{\tau}\right)\Bigg)\Bigg]$
@@
$\lambda$ denotes a scaling coefficient. The proposed regularizers improve semantic–spatial alignment in the latent space (Figure 3c) and reduce spurious correlations (Figure 9c). The learned attention maps can be leveraged for localized editing through attention-guided manipulation. Following  Ju et al. (2024), interventions are applied only to cross-attention weights corresponding to targeted tokens, while attention for non-intervened tokens is preserved. This ensures that edits remain spatially localized (e.g., gender) without altering unrelated attributes (e.g., hairstyle), improving identity preservation without introducing extra training objectives. Full algorithm is presented in Appendix C.3.